OpenStudy (anonymous):

Solve |3x - 4| = 15 A) {-19/3, 19/3} B) {11/3, 19/3} C) {-11/3, 19/3}

OpenStudy (gorv):

\[3x-4 = \pm 15\]

OpenStudy (gorv):

3x-4=15........3x=15+4......x=19/3 3x-4=-15......3x=-15+4......x=-11/3

OpenStudy (gorv):

ans = c